# Measuring Faithfulness of Chains of Thought by Unlearning Reasoning Steps

### v1

@@ 6 Results
In this section, we first report results of control measurements validating our intervention (§6.1). Subsequently, we report instance- and step-level faithfulness demonstrating the effectiveness of fur6.2). We then showcase one intended use of fur by identifying key reasoning steps and assess their plausibility through a user study (§6.3). Finally, we take a closer look at unlearning a single reasoning step and verify that unlearning has a deep effect on model parameters (§6.4).
Figure 3: A sample result of unlearning applied to a CoT step generated by LLaMA-3-8B on an instance from OpenbookQA. The bar charts represent no-CoT probability assigned to each answer option in that unlearning iteration. Model CoTs pre- and post-unlearning are displayed below. We omit CoTs from other unlearning iterations for space as they change very little after the 2nd iteration. Two steps are slightly shortened for presentation purposes.
@@ 6.1 Effectiveness of Unlearning
We report the results of unlearning when using the best hyperparameters for each method and dataset in Table 1. We measure each model’s efficacy, specificity, and MMLU performance before and after unlearning. The specificity and general capabilities of these models are largely unchanged while reporting good efficacy, indicating that the information from the target CoT step has been unlearned without affecting the model adversely. We report the results of various learning rates and discuss methodological choices in Section C.1.
@@ 6.2 Does Unlearning Change Predictions?
In the previous section, we show that we can unlearn information encoded in a reasoning step from the model. We now focus on how frequently unlearning information from reasoning steps causes the model predictions to change through ff-hard (Eq. 4), and contrast our method to Add-mistake, a contextual faithfulness method (Lanham et al., 2023). Add-mistake prompts another LM (in our case, gpt-4o-mini-2024-07-18) to introduce a mistake to a single step of a CoT generated by the target model. The target model is then prompted with a perturbed CoT containing the mistake. If the prediction of the model changes, the CoT is considered faithful. We report the results of instance-level faithfulness for fur and Add-mistake in Table 2.
@@
Models frequently change their prediction after unlearning. However, it is not clear how faithfulness relates to efficacy, i.e., if successful unlearning of a reasoning step is indicative of a change in prediction. We compute Pearson correlation between efficacy and ff-hard and observe a strong average correlation of $0.935$ with $p<0.0001$. We interpret this as indication that reasoning chains generated by the models are generally faithful, as the stronger we unlearn, the more frequent the change in prediction. The limiting factor is the interplay between efficacy and general capabilities, as stronger unlearning damages model integrity. Nevertheless, development of more precise unlearning techniques will remove this limitation. We discuss this further, along with step-level faithfulness Appendix F.
@@ 6.3 Quantifying Step Level Faithfulness
To quantitatively assess whether ff-soft identifies plausible steps as relevant, we conduct a user study on a random sample of $100$ instances. [^6] We randomly select instances from three bins of ff-soft depending on the amount and sign of mass moved from the initial prediction. See Appendix G for details. We show each participant a question, answer choices, and CoT steps, highlighting the answer predicted by the model and the target CoT step. We prompt the participants to annotate whether the step in question supports the predicted answer in context of the given CoT on a 1–5 Likert scale (Likert, 1932). We provide more details of the user study, data selection and the protocol in Appendix G.
Our results exhibit a weak Pearson correlation of $0.15$ between ff-soft and human ratings of supportiveness. If we filter out CoT steps where unlearning increases the probability of the initial answer, which are often cases where the model is uncertain of the prediction, the correlation increases to $0.27$, $p<0.02$. This result provides further evidence for findings from previous works showing that faithfulness, in general, does not correlate with plausibility (Agarwal et al., 2024). In order to improve correspondence between these two notions, one might need to specifically align LMs for reasoning plausibility (Ouyang et al., 2022).
Model Arc-Ch Book Sports SQA
LLaMA-8B $81.51$ $80.15$ $73.08$ $66.67$
LLaMA-3B $85.40$ $69.32$ $81.00$ $94.16$
Mistral-2 $83.87$ $90.50$ $80.34$ $86.49$
Phi-3 $75.74$ $75.54$ $69.23$ $73.58$
Table 3: LLM-as-judge results assessing if CoTs support different answers after unlearning. The percentage reported is how frequently GPT-4o states that the CoT supports a different answer post-unlearning.
6.4 Unlearning a Single Reasoning Step
To quantitatively assess how frequently the verbalized reasoning of the model changes post-unlearning, we use the LLM-as-a-judge paradigm (Zheng et al., 2023a) and verify if unlearning has caused the verbalized reasoning to support a different answer, indicating deeper unlearning, or if the change in model prediction is caused by shallow unlearning, which does not affect the reasoning of the model Cohen et al. (2024). We first select instances where both CoT and no-CoT models agree in their changed predictions. From these cases, we select reasoning steps from the last iteration of unlearning. We prompt gpt-4o-mini-2024-07-18 to judge whether the CoTs generated by the model before and after unlearning support different answers. We report the results in Table 3 and detail the prompting setup in Appendix E.
@@ Acknowledgments
Appendix A Dataset and Model Statistics
@@
We report the base performance of the analyzed models on the datasets we selected, with and without CoT in Table 4. Statistics on the total, and average counts of CoT steps can be seen in Table 5. We describe and exemplify the prompting setup in Appendix B.
@@
Model CoT Arc-Ch Book Sports SQA
LLaMA-8B $0.817$ $0.704$ $0.822$ $0.679$
$0.839$ $0.778$ $0.836$ $0.736$
LLaMA-3B $0.726$ $0.674$ $0.500$ $0.609$
$0.774$ $0.757$ $0.561$ $0.652$
Mistral-2 $0.709$ $0.739$ $0.711$ $0.625$
$0.774$ $0.730$ $0.719$ $0.701$
Phi-3 $0.909$ $0.804$ $0.610$ $0.622$
$0.870$ $0.848$ $0.789$ $0.713$
Table 4: Results of analyzed models on the datasets when promted with and without CoTs.
Dataset # CoT steps Avg. # steps # Inst
ARC-Challenge $1803$ $7.84$ $230$
OpenBookQA $1830$ $7.96$ $230$
Sports $1415$ $6.21$ $228$
StrategyQA $1956$ $8.50$ $230$
Table 5: Statistics of analyzed datasets in terms of instances and CoT steps.
Appendix B MCQA Task Prompts
@@
Appendix C Unlearning Setup & Hyperparameters
We adapt the implementation of NPO-KL from the official repository. [^7] https://github.com/licong-lin/negative-preference-optimization We use the best hyperparameters found by the original paper (Zhang et al., 2024b) except for the values which we highlight in bold. See Table 6 for values.
@@ Appendix C Unlearning Setup & Hyperparameters
Table 6: Hyperparameters used in the implementation of NPO-KL. Bold values deviate from the original paper.
@@
When performing unlearning, we backpropagate only on target tokens which are constituents of content words, namely nouns, proper nouns, adjectives, verbs and numbers. We filter out and don’t unlearn all CoT steps which do not have at least two target tokens. This usually corresponds to the index in the CoT step enumeration which plenty of models produce (e.g. 1. This is a CoT step), which is sentencized as a standalone sentence by SpaCy.
When unlearning, NPO-KL uses KL regularization to control updates to model parameters, which could otherwise be unbounded (Zhang et al., 2024b). During optimization, the model is regularized not to deviate from its initial version with respect to KL divergence of the predictive distribution on a retain set. For the retain set, we select a random sample of $4$ other CoT steps from the same dataset. We perform the same filtering in the retain set, keeping only steps which contain more than two tokens which are constituents of content words, and only target those words for KL regularization.
C.1 Learning Rate Selection
@@
Figure 5: Learning rate selection results for NPO-KL. Experiments ran on $30$ instances for all datasets. Size of the marker depicts faithfulness, only for information purposes — faithfulness was not used as the selection criterion. Learning rates omitted for clarity, but as a rule, the higher the learning rate, the higher efficacy, and the lower the specificity. Figure presented for glance-value, scores are also reported in tabular format in Table 7.
Arc-Challenge OpenbookQA Sports StrategyQA
Model LR Eff Spec FF Eff Spec FF Eff Spec FF Eff Spec FF
LLaMA-8B $1\mathrm{e}{-06}$ $0.4$ $99.2$ $6.7$ $0.6$ $97.4$ $3.3$ $0.7$ $98.5$ $10.0$- -- --
$3\mathrm{e}{-06}$ $3.3$ $99.1$ $13.3$ $4.4$ $97.5$ $6.7$ $6.1$ $98.7$ $13.3$ $4.6$ $99.2$ $6.7$
$5\mathrm{e}{-06}$ $13.1$ $98.9$ $20.0$ $15.2$ $97.5$ $16.7$ $20.7$ $98.1$ $26.7$ $16.0$ $98.2$ $10.0$
$1\mathrm{e}{-05}$ $35.2$ $97.6$ $46.7$ $37.0$ $97.2$ $43.3$ $44.9$ $94.0$ $43.3$ $39.4$ $94.8$ $33.3$
$3\mathrm{e}{-05}$ $66.0$ $91.2$ $60.0$ $68.0$ $87.6$ $73.3$- -- -- $69.5$ $78.9$ $86.7$
$5\mathrm{e}{-05}$ $75.7$ $81.2$ $70.0$- -- -- $77.6$ $57.8$ $80.0$ $77.0$ $67.4$ $90.0$
$0.0001$- -- -- -- -- -- -- -- -- $80.6$ $59.5$ $96.7$
LLaMA-3B $5\mathrm{e}{-06}$ $1.6$ $97.0$ $10.0$- -- -- $1.4$ $100.0$ $3.3$ $2.0$ $100.0$ $13.3$
$1\mathrm{e}{-05}$ $6.5$ $97.7$ $30.0$ $7.9$ $99.3$ $23.3$ $5.3$ $100.0$ $13.3$ $7.7$ $99.9$ $23.3$
$3\mathrm{e}{-05}$ $31.3$ $97.4$ $76.7$ $36.0$ $94.8$ $60.0$ $27.6$ $96.4$ $53.3$ $34.5$ $96.7$ $70.0$
$5\mathrm{e}{-05}$- -- -- $56.8$ $90.4$ $90.0$ $49.4$ $85.9$ $80.0$ $56.3$ $87.7$ $83.3$
$0.0001$ $69.3$ $81.2$ $96.7$ $73.0$ $70.7$ $96.7$ $68.9$ $80.2$ $86.7$ $73.3$ $66.3$ $96.7$
Mistral-2 $1\mathrm{e}{-06}$ $11.4$ $100.0$ $10.0$ $12.5$ $100.0$ $13.3$- -- -- -- -- --
$3\mathrm{e}{-06}$ $43.6$ $99.0$ $30.0$ $43.6$ $99.2$ $33.3$ $43.7$ $93.2$ $40.0$ $41.7$ $97.2$ $33.3$
$5\mathrm{e}{-06}$ $60.8$ $95.6$ $46.7$ $60.2$ $96.7$ $56.7$ $60.3$ $85.4$ $60.0$ $58.7$ $94.9$ $53.3$
$1\mathrm{e}{-05}$ $74.1$ $89.1$ $73.3$ $73.6$ $91.4$ $73.3$ $73.6$ $71.5$ $70.0$ $72.7$ $86.3$ $76.7$
$3\mathrm{e}{-05}$ $80.6$ $75.5$ $96.7$ $80.1$ $64.9$ $80.0$- -- -- -- -- --
Phi-3 $3\mathrm{e}{-05}$ $3.6$ $100.0$ $6.7$ $4.0$ $100.0$ $16.7$ $8.0$ $97.9$ $30.0$ $4.4$ $99.8$ $10.0$
$5\mathrm{e}{-05}$- -- -- $13.2$ $100.0$ $23.3$ $25.1$ $96.8$ $50.0$ $13.8$ $97.6$ $16.7$
$0.0001$ $34.4$ $99.4$ $53.3$ $38.5$ $99.4$ $46.7$ $55.8$ $90.9$ $66.7$ $39.6$ $92.8$ $53.3$
$0.0003$ $69.2$ $93.7$ $76.7$ $70.7$ $92.6$ $76.7$- -- -- -- -- --
@@
$0.001$ $80.7$ $59.1$ $96.7$ $80.8$ $49.1$ $93.3$- -- -- -- -- --
Table 7: Learning rate selection results for NPO-KL. Experiments ran on $30$ instances for all datasets. Faithfulness was not used as the selection criterion, but is here only for informativeness. Best learning rates per model & dataset highlighted. Criterion was $\max(\text{efficacy})\hskip 5.0pt\text{s.t.}\hskip 5.0ptround(\text{specificity})\geq 95$.
Appendix D Add-mistake Implementation

### v2

@@ 6 Results
We first report results of control measurements validating our intervention (§6.1). Subsequently, we demonstrate the effectiveness of fur in detecting faithful reasoning (§6.2). We then investigate the effect of unlearning on generated CoTs (§6.3). Finally, we use fur to identify key reasoning steps and assess their plausibility in a user study (§6.4).
@@ 6.1 Effectiveness of Unlearning
We report the results of unlearning when using the best hyperparameters for each method and dataset in Table 1. We measure each model’s efficacy, specificity, and MMLU performance before and after unlearning. The specificity and general capabilities of these models are largely unchanged while reporting good efficacy, indicating that the information from the target CoT step has been unlearned without affecting the model adversely. We report the results of various learning rates and discuss methodological choices in Appendix E.
@@ 6.2 Does Unlearning Change Predictions?
In the previous section, we show that we can precisely unlearn information encoded in a reasoning step. We now focus on how frequently unlearning information from reasoning steps causes the model predictions to change through ff-hard (Eq. 4), and contrast our method to Add-mistake, a contextual faithfulness method (Lanham et al., 2023). Add-mistake prompts another LM (in our case, gpt-4o-mini-2024-07-18) to introduce a mistake to a single step of a CoT generated by the target model. The target model is then prompted with a perturbed CoT containing the mistake. If the prediction of the model changes, the CoT is considered faithful. We report the results of instance-level faithfulness for fur and Add-mistake in Table 2.888We explore and comment other baselines in Appendix B.
@@
We find that unlearning efficacy is highly indicative of faithfulness. The Pearson correlation between average efficacy and ff-hard is high: $0.889$ with $p<0.0001$. We interpret this as indication that reasoning chains generated by the models are generally faithful, as the stronger we unlearn, the more frequent the change in prediction. The limiting factor is that stronger unlearning damages model integrity. Nevertheless, development of more precise unlearning techniques will remove this limitation. We discuss this further, along with step-level faithfulness Appendix H.
Figure 3: A sample result of unlearning applied to a CoT step generated by LLaMA-3-8B on an instance from OpenbookQA. The bar charts represent no-CoT probability assigned to each answer option in that unlearning iteration. Model CoTs pre- and post-unlearning are displayed below. We omit CoTs from other unlearning iterations for space as they change very little after the 2nd iteration. One step is slightly shortened for presentation purposes.
Model Arc-ch Book Sports SQA TQA
LLaMA-8B $81.5$ $80.2$ $73.1$ $66.7$ $86.9$
LLaMA-3B $85.4$ $69.3$ $81.0$ $94.2$ $84.9$
Mistral-2 $83.9$ $90.5$ $80.3$ $86.5$ $81.7$
Phi-3 $75.7$ $75.5$ $69.2$ $73.6$ $81.1$
Table 3: LLM-as-a-judge results assessing if CoTs support different answers after unlearning. Numbers are percentages of how frequently GPT-4o states that the CoT supports a different answer post-unlearning.
6.3 Does Unlearning Change Reasoning?
To quantitatively assess how frequently the verbalized reasoning of the model changes post-unlearning, we employ an LLM-as-a-judge (Zheng et al., 2023a) to verify if unlearning caused the generated CoT to support a different answer, indicating deeper unlearning, or if the change in model prediction is not reflected in reasoning Cohen et al. (2024). We first select instances where CoT and no-CoT models agree in their changed predictions. From these cases, we select reasoning steps from the last iteration of unlearning. We prompt gpt-4o-mini-2024-07-18 to judge whether the CoTs generated by the model before and after unlearning support different answers. We report results in Table 3 and detail our setup in Appendix G.
@@ 6.4 Quantifying Step Level Faithfulness
To quantitatively assess whether ff-soft identifies plausible steps as relevant, we conduct a user study on a random sample of $100$ instances. We show each participant a question, answer choices, and CoT steps, highlighting the answer predicted by the model and the target CoT step. We prompt the participants to annotate whether the step in question supports the predicted answer in context of the given CoT on a 1–5 Likert scale (Likert, 1932). We provide more details of the user study, data selection and the protocol in Appendix I.
We find a weak Pearson correlation of $0.15$ between ff-soft and human ratings of supportiveness. This result provides further evidence that faithfulness, in general, does not correlate with plausibility (Agarwal et al., 2024). In order to improve correspondence between these two notions, one might need to specifically align LMs for reasoning plausibility (Ouyang et al., 2022).
@@ Acknowledgments
Appendix A Background on Machine Unlearning
Motivated by the need to erase sensitive information from machine learning models (Cao and Yang, 2015; Harding et al., 2019; Ippolito et al., 2023), the field of machine unlearning emerged as an efficient alternative to filter-then-retrain-based approaches (Neel et al., 2021; Jang et al., 2023; Eldan and Russinovich, 2023; Liu et al., 2024, inter alia). Machine unlearning methods update parameters of the model in a way that reduces the competency of the model on unwanted data (henceforth, forget), while retaining general capabilities through regularization on retain data. Such methods decrease token probabilities on the forget data by gradient-based approaches (Jang et al., 2023; Eldan and Russinovich, 2023; Gandikota et al., 2024; Li et al., 2024; Zhang et al., 2024b) or directly updating parameters (Meng et al., 2022, 2023; Wu et al., 2023; Ashuach et al., 2024).
In order to ensure that unlearning does not adversely affect the model to the point it is unusable, LMs need to satisfy the following desiderata post-unlearning: (1) efficacy, controlling whether the forget data was erased from the model; (2) specificity, controlling that the edit is localized to the target information, often by probing the model on closely related data; (3) general capabilities, measuring whether the model retains fluency and performance on unrelated data.
Negative Preference Optimization.
In this work, we use a preference-optimization based unlearning method: negative preference optimization (NPO; Zhang et al., 2024b). The core idea underpinning NPO is rooted in gradient ascent. The initial model $\pi_{\mathcal{D}}$ is trained on a mixture of wanted and unwanted data $\mathcal{D}=\mathcal{D}_{FG}\cup\mathcal{D}_{RT}$. Our goal is to eliminate unwanted information $\mathcal{D}_{FG}$ from the model. Therefore, applying the reverse language modeling objective on the forget data:
$\mathcal{L}_{\text{GA}}(\theta)=\mathbb{E}_{\mathcal{D}_{\text{FG}}}[\text{log}(\pi_{\theta}(y|x))],$ (6)
would approximately revert the optimization, producing $\pi_{\mathcal{D}_{RT}}$.
Applying gradient ascent in this way runs into two practical issues. Firstly, we often do not have access to the training dataset, and therefore neither to the unwanted data $\mathcal{D}_{\text{FG}}$. Secondly, the gradient ascent objective is unbounded by virtue of maximizing the next-token prediction loss, frequently resulting in catastrophic collapse (Zhang et al., 2024b). Machine unlearning approaches resolve the first issue by approximating $\mathcal{D}_{\text{FG}}$ with a forget set containing samples of unwanted data. NPO resolves the second issue by constraining that the policy (predictive distribution) of the unlearned model $\pi_{\theta}$ should not diverge too far from a reference model $\pi_{\text{ref}}$. In practice, the frozen base model is used as the reference ($\pi_{\text{ref}}=\pi_{\mathcal{D}}$).
The NPO loss is then defined as:
$\mathcal{L}_{\text{NPO},\beta}(\theta)=\frac{2}{\beta}\mathbb{E}_{\mathcal{D}_{\text{FG}}}\left[\text{log}\left(1+\left(\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}\right)^{\beta}\right)\right],$ (7)
where $\beta>0$ is the inverse temperature. This loss can be viewed as Direct Preference Optimization (DPO; Rafailov et al., 2023) without the positive samples. In practice, along with the loss term in Eq. (7), NPO also constrains the KL divergence between the unlearned and reference models on retain data in order to guarantee fluency, similar to other works (Li et al., 2024; Gandikota et al., 2024):
$\mathcal{K}_{\text{RT}}=\mathbb{E}_{\mathcal{D}_{\text{RT}}}\left[D_{\text{KL}}\big{(}\pi_{\theta}(\cdot|x)\,\middle\|\,\pi_{\text{ref}}(\cdot|x)\big{)}\right].$ (8)
As the KL divergence regularizer maintains that the unlearned model does not diverge from the reference, it does not introduce new information to the model.
In our work, the forget set consists of all tokens of a given content word from a CoT step as the output $y$ to be forgotten, paired with the word’s preceding context as the input $x$. The retain data is constructed similarly out of CoT steps from unrelated instances. We alter a subset $\theta$ of the full model’s parameters by minimizing:
$\mathcal{L}=\mathcal{L}_{\text{NPO},\beta}(\theta)+\mathcal{K}_{\text{RT}}.$ (9)
Appendix B Alternatives to Measuring (Un)Faithfulness
A number of methods have been proposed with the goal of measuring faithfulness of model reasoning (Lanham et al., 2023; Atanasova et al., 2023; Bentham et al., 2024; Siegel et al., 2024; Chen et al., 2024; Madsen et al., 2024, inter alia). However, not all of them are applicable to our setup. Some of the methods are designed for NLI tasks (Atanasova et al., 2023; Siegel et al., 2024; Parcalabescu and Frank, 2024) while others aim at multi-hop QA tasks (Parcalabescu and Frank, 2024; Lanham et al., 2023; Chen et al., 2024; Bentham et al., 2024). In our work, we focus on QA tasks as datasets requiring multi-hop reasoning are more prominent in this task (Jacovi et al., 2024), allowing us a broader domain coverage. We further choose for MCQA, as alternative answers in these datasets are often by design plausible, and this allows for a more in-depth analysis of how unlearning affects the underlying reasoning of the model, by e.g. making it opt for plausible alternatives. On the contrary, in NLI, the model is either right or wrong – there are few “alternative explanations” and the analysis one can do is limited.
Other Baselines.
We have considered other baselines applicable to CoT reasoning (cf. Table 1. in Zaman and Srivastava, 2025). Namely, we explore Early Answering, Filler Tokens, Adding Mistakes and Paraphrasing (Lanham et al., 2023) as well as CC-SHAP (Parcalabescu and Frank, 2024). Adding Mistakes is a simple contextual faithfulness method that works well, and we compare to its results in Table 2. We replicate the Paraphrasing setup and find that it is able to identify a small proportion ($2.84$%, on average) of instances as unfaithful. See full results in Table 4. Importantly, this does not imply that the remaining instances are faithful, which is the goal of our work.
Model Arc-Ch Book Sports SQA
LLaMA-8B $2.60$ $1.55$ $1.72$ $2.15$
LLaMA-3B $3.39$ $4.65$ $1.19$ $0.57$
Mistral-2 $9.73$ $3.70$ $3.68$ $3.12$
Phi-3 $3.81$ $2.56$ $0.00$ $2.65$
Table 4: The percentage of CoTs identified as unfaithful by the Paraphrase baseline (Lanham et al., 2023).
The remaining methods from Lanham et al. (2023) aim to identify whether reasoning is produced post-hoc, or truly necessary to produce the prediction. Early Answering truncates the CoT, while Filler Tokens substitutes the CoT with ellipsis tokens. Then, if the answer did not change, the CoT is deemed unnecessary (post-hoc reasoning). These measures do not aim to determine faithfulness of CoTs. Post-hoc reasoning can still be a true verbalization of latent reasoning, which is what both fur and Add-Mistake find Table 2 since we only evaluate faithfulness cases where CoT and no-CoT predictions of models agree. Such cases would be identified as post-hoc reasoning according to Early Answering and Filler Tokens, but should not be discarded as unfaithful.
Finally, we experiment with CC-SHAP (Parcalabescu and Frank, 2024), a self-consistency measure based on Shapley values, which measures the convergence between input tokens salient for the prediction and explanation. We use the official implementation from the authors [^9] https://github.com/Heidelberg-NLP/CC-SHAP, but when applying SHAP to instances from our dataset, relative importances of tokens from input are frequently exactly zero (importances for reasoning do not behave in this manner), which results in NaN CC-SHAP scores in $90.24$% of instances across datasets and models. We believe such low scores for model predictions are caused by the fact that the inputs are only the question and answer options, while the evidence (reasoning) is intrinsic to the model.
Appendix C Dataset and Model Statistics
@@
We report the base performance of the analyzed models on the datasets we selected, with and without CoT in Table 5. Statistics on the total, and average counts of CoT steps can be seen in Table 6. We describe and exemplify the prompting setup in Appendix D.
@@
Model CoT Arc-Ch Book Sports SQA TQA
LLaMA-8B $0.82$ $0.70$ $0.82$ $0.68$ $0.44$
$\underline{0.84}$ $\underline{0.78}$ $\underline{0.84}$ $\underline{0.74}$ $\underline{0.52}$
LLaMA-3B $0.73$ $0.67$ $0.50$ $0.61$ $0.55$
$\underline{0.77}$ $\underline{0.76}$ $\underline{0.56}$ $\underline{0.65}$ $\underline{0.57}$
Mistral-2 $0.71$ $0.74$ $0.71$ $0.63$ $0.35$
$\underline{0.77}$ $0.73$ $0.72$ $\underline{0.70}$ $\underline{0.46}$
Phi-3 $\underline{0.91}$ $0.80$ $0.61$ $0.62$ $0.59$
$0.87$ $\underline{0.85}$ $\underline{0.79}$ $\underline{0.71}$ $0.59$
Table 5: Results of analyzed models on the datasets when prompted with and without CoTs. Results better by at least one percentage point underlined. In general, the tasks are difficult for the models, and using CoT improves over no-CoT.
Model Arc-Ch Book Sports SQA TQA
LLaMA-8B $4.36$ $4.24$ $3.96$ $3.90$ $5.52$
LLaMA-3B $7.25$ $6.71$ $7.29$ $8.45$ $7.34$
Mistral-2 $3.65$ $3.70$ $4.85$ $4.55$ $4.11$
Phi-3 $7.75$ $7.91$ $6.20$ $8.46$ $10.20$
Table 6: Average number of CoT steps per model and dataset, measured on the full $250$ instances from each dataset ($248$ for Sports).
Appendix D MCQA Task Prompts
@@
Appendix E Unlearning Setup & Hyperparameters
We adapt the implementation of NPO+KL from the official repository. [^10] https://github.com/licong-lin/negative-preference-optimization We use the best hyperparameters found by the original paper (Zhang et al., 2024b) except for the values which we highlight in bold. See Table 7 for values.
@@ Appendix E Unlearning Setup & Hyperparameters
Table 7: Hyperparameters used in the implementation of NPO+KL. Bold values deviate from the original paper.
@@
When performing unlearning, we backpropagate only on target tokens which are constituents of content words, namely nouns, proper nouns, adjectives, verbs and numbers. We filter out and don’t unlearn all CoT steps which do not have at least two target tokens. This usually corresponds to the index in the CoT step enumeration which plenty of models produce (e.g. 1. This is a CoT step), where “1.” is sentencized as a standalone sentence by SpaCy.
When unlearning, NPO+KL uses KL regularization to control updates to model parameters, which could otherwise be unbounded (Zhang et al., 2024b). During optimization, the model is regularized not to deviate from its initial version with respect to KL divergence of the predictive distribution on a retain set. For the retain set, we select a random sample of $4$ other CoT steps from the same dataset. We perform the same filtering in the retain set, keeping only steps which contain more than two tokens which are constituents of content words, and only target those words for KL regularization.
E.1 Learning Rate Selection
@@
Figure 5: Learning rate selection results for NPO+KL. Experiments ran on $30$ instances for all datasets. Size of the marker depicts faithfulness, only for information purposes — faithfulness was not used as the selection criterion. Learning rates omitted for clarity, but as a rule, the higher the learning rate, the higher the efficacy, and the lower the specificity. Figure presented for glance-value, scores are also reported in tabular format in Table 8.
Arc-Challenge OpenbookQA Sports StrategyQA TruthfulQA
LR Eff Spec FF Eff Spec FF Eff Spec FF Eff Spec FF Eff Spec FF
LLaMA-8B $1\mathrm{e}{-06}$ $0.4$ $99.2$ $6.7$ $0.6$ $97.4$ $3.3$ $0.7$ $98.5$ $10.0$- -- -- -- -- --
$3\mathrm{e}{-06}$ $3.3$ $99.1$ $13.3$ $4.4$ $97.5$ $6.7$ $6.1$ $98.7$ $13.3$ $4.6$ $99.2$ $6.7$- -- --
$5\mathrm{e}{-06}$ $13.1$ $98.9$ $20.0$ $15.2$ $97.5$ $16.7$ $20.7$ $98.1$ $26.7$ $16.0$ $98.2$ $10.0$ $15.8$ $98.8$ $43.3$
$1\mathrm{e}{-05}$ $35.2$ $97.6$ $46.7$ $37.0$ $97.2$ $43.3$ $44.9$ $94.0$ $43.3$ $39.4$ $94.8$ $33.3$ $39.2$ $97.2$ $60.0$
$3\mathrm{e}{-05}$ $66.0$ $91.2$ $60.0$ $68.0$ $87.6$ $73.3$- -- -- $69.5$ $78.9$ $86.7$ $69.8$ $82.0$ $86.7$
$5\mathrm{e}{-05}$ $75.7$ $81.2$ $70.0$- -- -- $77.6$ $57.8$ $80.0$ $77.0$ $67.4$ $90.0$ $77.5$ $62.9$ $90.0$
$0.0001$- -- -- -- -- -- -- -- -- $80.6$ $59.5$ $96.7$- -- --
LLaMA-3B $5\mathrm{e}{-06}$ $1.6$ $97.0$ $10.0$- -- -- $1.4$ $100.0$ $3.3$ $2.0$ $100.0$ $13.3$ $1.0$ $99.9$ $16.7$
$1\mathrm{e}{-05}$ $6.5$ $97.7$ $30.0$ $7.9$ $99.3$ $23.3$ $5.3$ $100.0$ $13.3$ $7.7$ $99.9$ $23.3$ $4.3$ $99.5$ $46.7$
$3\mathrm{e}{-05}$ $31.3$ $97.4$ $76.7$ $36.0$ $94.8$ $60.0$ $27.6$ $96.4$ $53.3$ $34.5$ $96.7$ $70.0$ $24.9$ $95.3$ $83.3$
$5\mathrm{e}{-05}$- -- -- $56.8$ $90.4$ $90.0$ $49.4$ $85.9$ $80.0$ $56.3$ $87.7$ $83.3$ $47.8$ $85.4$ $100.0$
$0.0001$ $69.3$ $81.2$ $96.7$ $73.0$ $70.7$ $96.7$ $68.9$ $80.2$ $86.7$ $73.3$ $66.3$ $96.7$ $69.4$ $63.0$ $100.0$
Mistral-2 $1\mathrm{e}{-06}$ $11.4$ $100.0$ $10.0$ $12.5$ $100.0$ $13.3$- -- -- -- -- -- $10.8$ $99.1$ $30.0$
$3\mathrm{e}{-06}$ $43.6$ $99.0$ $30.0$ $43.6$ $99.2$ $33.3$ $43.7$ $93.2$ $40.0$ $41.7$ $97.2$ $33.3$ $40.8$ $94.6$ $63.3$
$5\mathrm{e}{-06}$ $60.8$ $95.6$ $46.7$ $60.2$ $96.7$ $56.7$ $60.3$ $85.4$ $60.0$ $58.7$ $94.9$ $53.3$ $57.4$ $84.9$ $83.3$
$1\mathrm{e}{-05}$ $74.1$ $89.1$ $73.3$ $73.6$ $91.4$ $73.3$ $73.6$ $71.5$ $70.0$ $72.7$ $86.3$ $76.7$ $72.4$ $68.3$ $90.0$
$3\mathrm{e}{-05}$ $80.6$ $75.5$ $96.7$ $80.1$ $64.9$ $80.0$- -- -- -- -- -- $80.2$ $41.2$ $93.3$
Phi-3 $3\mathrm{e}{-05}$ $3.6$ $100.0$ $6.7$ $4.0$ $100.0$ $16.7$ $8.0$ $97.9$ $30.0$ $4.4$ $99.8$ $10.0$ $2.5$ $98.8$ $13.3$
$5\mathrm{e}{-05}$- -- -- $13.2$ $100.0$ $23.3$ $25.1$ $96.8$ $50.0$ $13.8$ $97.6$ $16.7$ $8.8$ $96.7$ $46.7$
$0.0001$ $34.4$ $99.4$ $53.3$ $38.5$ $99.4$ $46.7$ $55.8$ $90.9$ $66.7$ $39.6$ $92.8$ $53.3$ $29.1$ $90.2$ $83.3$
$0.0003$ $69.2$ $93.7$ $76.7$ $70.7$ $92.6$ $76.7$- -- -- -- -- -- -- -- --
@@
$0.001$ $80.7$ $59.1$ $96.7$ $80.8$ $49.1$ $93.3$- -- -- -- -- -- $73.6$ $59.0$ $96.7$
Table 8: Learning rate selection results for NPO+KL. Experiments ran on $30$ instances for all datasets. Faithfulness was not used as the selection criterion, but is here only for informativeness. Best learning rates per model & dataset highlighted. Criterion was $\max(\text{efficacy})\hskip 5.0pt\text{s.t.}\hskip 5.0ptround(\text{specificity})\geq 95$.
Appendix F Add-mistake Implementation